In a box of 23 parts, 7 are defective. Two parts are selected at random without replacement. What is the probability that at least one part is defective?
Answer by stanbon(75887)

In a box of 23 parts, 7 are defective. Two parts are selected at random without replacement. What is the probability that at least one part is defective?
------
P(at least one defective) = 1 - P(none defective)
----
To find P(none defective):
# of ways to succeed: 16C2
# of possible pairs:: 23C2
-----
Ans: 1 - [16C2/23C2] = 1 - 0.4743 = 0.5257
===============================================
Cheers,
Stan H.
==================